OpenStudy (anonymous):

which of the following is between 10 and 11..

OpenStudy (anonymous):

what are the options ?

OpenStudy (anonymous):

\[\sqrt{83} \sqrt{118} \sqrt{126} \sqrt{132}\]

OpenStudy (anonymous):

\[\sqrt{118}\]

OpenStudy (anonymous):

thank you ...

OpenStudy (anonymous):

your welcome :]

OpenStudy (anonymous):

how you got that answer

OpenStudy (anonymous):

go to the calculator and type in 118 and find the square root and its 10. something

OpenStudy (anonymous):

okay thanks